OpenStudy (anonymous):

what are the endpoints of major axis (3,2) and (3,-4) and endpoints of minor axis(5,-1) and(1,-1)

OpenStudy (anonymous):

you want equation of ellipse?

OpenStudy (anonymous):

i have and equation but my answer keep coming out wrong

OpenStudy (anonymous):

what did you the for lenght of major axis and minor axis

OpenStudy (anonymous):

These are different numbers the first two sets i got wrong so i sought out help

OpenStudy (anonymous):

Major axis I got 2-(-4))=6 Minor axis I got 5-1=4

OpenStudy (anonymous):

x^2/a^2 +y^2/b^2 =1